### v1

# BihoT: A Large-Scale Dataset and Benchmark for Hyperspectral Camouflaged Object Tracking

Abstract
Hyperspectral object tracking (HOT) has exhibited potential in various applications, particularly in scenes where objects are camouflaged. Existing trackers can effectively retrieve objects via band regrouping because of the bias in existing HOT datasets, where most objects tend to have distinguishing visual appearances rather than spectral characteristics. This bias allows the tracker to directly use the visual features obtained from the false-color images generated by hyperspectral images without the need to extract spectral features. To tackle this bias, we find that the tracker should focus on the spectral information when object appearance is unreliable. Thus, we provide a new task called hyperspectral camouflaged object tracking (HCOT) and meticulously construct a large-scale HCOT dataset, termed BihoT, which consists of 41,912 hyperspectral images covering 49 video sequences. The dataset covers various artificial camouflage scenes where objects have similar appearances, diverse spectrums, and frequent occlusion, making it a very challenging dataset for HCOT. Besides, a simple but effective baseline model, named spectral prompt-based distractor-aware network (SPDAN), is proposed, comprising a spectral embedding network (SEN), a spectral prompt-based backbone network (SPBN), and a distractor-aware module (DAM). Specifically, the SEN extracts spectral-spatial features via 3-D and 2-D convolutions to form a refined prompt representation. Then, the SPBN fine-tunes powerful RGB trackers with spectral prompts and alleviates the insufficiency of training samples. Moreover, the DAM utilizes a novel statistic to capture the distractor caused by occlusion from objects and background and corrects the deterioration of the tracking performance via a novel motion predictor. Extensive experiments demonstrate that our proposed SPDAN achieves state-of-the-art performance on the proposed BihoT and other HOT datasets.
Index Terms:
Hyperspectral camouflaged object tracking, prompt-based learning, HCOT dataset, feature fusion.I Introduction
Visual object tracking [1, 2] is a crucial task that involves finding the same objects in subsequent frames based on their initial positions in the first frame. It has great potential in various applications such as autonomous driving, video analysis, and so on [3, 4, 5]. However, visual object tracking makes it difficult to handle interference caused by similar appearances and background clutter. Hyperspectral images (HSIs) contain the texture and spectral information of an object and can use spectral information to track an object in the presence of visual interference. Hence, hyperspectral object tracking (HOT) has attracted more attention from the research community.
With the construction of HOT datasets such as HOT-2020 [6], IMEC-25 [7], and WHU-Hi-H3 [8], many deep learning-based methods have been proposed [9, 10, 11, 12, 13]. Although the above methods have achieved great progress, many discriminative spectral features still have not been explored. As shown in Fig. 1(a), we argue that the existing HOT dataset has two characteristics: one is the visual distinguishable (v-dis) factor, which means that using visual information can correctly track objects, and the other is the spectral distinguishable (s-dis) factor, which means that using spectral information can correctly track objects. Taking the HOTC-2020 dataset as an example, video sequences with the v-dis factor account for 77.1% of the entire dataset, while the s-dis factor only accounts for 22.9% of the entire dataset. Considering that most existing HS tracking methods rely on pre-trained RGB trackers for processing false-color images generated from HS images [10, 11, 14], which are good at handling v-dis factor. So in many cases, even if the tracker cannot extract robust spectral features, it can still capture objects through visual information (v-dis factor). This bias of datasets makes most algorithms overlook the utilization of the advantages of hyperspectral information.
Despite existing HOT tasks exploring the degradation of object visual information due to factors like illumination changes and motion blur, objects are typically in a v-dis state. This condition does not make the trackers leverage the spectral characteristics of objects, thereby limiting their ability to handle tracking tasks in complex environments. To address this problem, we introduce a novel hyperspectral camouflaged object tracking (HCOT) task, which focuses on tracking objects that share the same visual information but differ in spectral information. Considering that no existing dataset suits this task, we construct the first large-scale HCOT dataset, BihoT. This dataset has two significant advantages over existing datasets. First, BihoT collects many visually similar objects, called camouflaged objects. As illustrated in Fig. 1(b), the plastic fake fruit is a camouflaged object relative to the real fruit. The proportion of the s-dis factor reaches 95.5%, while the proportion of the v-dis factor is only 4.5%. For camouflaged objects, the v-dis factor is usually unreliable in tracking, underscoring the importance of the spectral (s-dis) factor. Second, BihoT consists of 49 hyperspectral video (HSV) sequences, containing 41,912 HSI in total, making it the largest HCOT dataset with 25 bands. Each HS image encompasses 25 bands and a resolution of 409 217, meeting the data requirements for deep learning.
Furthermore, to promote the development of HCOT, we propose a spectral prompt-based distractor-aware network (SPDAN) as a simple yet effective baseline to tackle the challenges of spectral feature extraction and frequent occlusion. Specifically, a spectral embedding network (SEN) is designed to transform the spatial and spectral information to the tokenized features, which offers discriminative spectral information for existing patch embedding methods [15] via a 3-D convolutional layer. Then, motivated by prompt learning [16], a spectral prompt-based backbone network (SPBN) is proposed to transfer the knowledge from the RGB domain to the HS domain to extract the robust spectral-aware features. Finally, to address the challenges of frequent occlusion in practical applications, we propose a distractor-aware module (DAM) to capture frames with disturbances such as occlusion accurately. Different from the SENSE [10], DAM considers both APCE [17] and the offset of the prediction boxes across frames, which can make more accurate judgments on distractors and adopt motion state models to predict states with occlusion, effectively mitigating the decline in the tracking performance caused by such distractors. As shown in Fig. 2, it can be observed that our proposed method outperforms the other state-of-the-art HOT methods on the HOTC-2020 dataset. The main contributions of this paper are as follows:
1.
The first large-scale HCOT dataset, BihoT, is constructed, which contains 41,912 HS camouflaged images paired with the bounding box annotations. It provides more challenging scenes and brings profound insights for HOT.
2.
A novel spectral embedding network is designed to transfer the spatial and spectral information to the tokenized features, and aggregate the spectral and visual features from the prompt-based learning to strengthen camouflaged feature representation.
3.
A distractor-aware module is proposed to perceive the distractor by considering both the average peak-to-correlation energy and the offset of the predicted boxes, and rectify the tracking results through a motion predictor, improving the robustness of the tracker to distractors.
The remaining structure of the paper is structured as follows. Section II reviews some related works. Section III introduces the proposed BihoT dataset. Section IV formulated the proposed methods. Section V provides the ablation and comparative experimental results. Finally, conclusions are drawn in Section VI.
| Datasets | Year | Band | Training Num. | Testing Num. | Labeled Sequence Num. | Avg. #images/sequences | Annotations |
| HOTC-2020[6] | 2020 | 16 | 12,839 | 16,574 | 75 | 392.17 | 29,413 |
| IMEC-25[7] | 2021 | 25 | 14,062 | 14,357 | 135 | 210.51 | 28,419 |
| WHU-Hi-H3[8] | 2022 | 25 | 22,753 | 6969 | 9 | 774.33 | 6969 |
| HOTC-2023-RedNIR[18] | 2023 | 15 | 5737 | 3500 | 14 | 659.78 | 9237 |
| HOTC-2023-VIS[18] | 2023 | 16 | 19,228 | 20,074 | 55 | 714.58 | 39,302 |
| HOTC-2023-NIR[18] | 2023 | 25 | 8161 | 4856 | 40 | 325.42 | 13,017 |
| BihoT | 2024 | 25 | 22,992 | 18,920 | 49 | 855.35 | 41,912 |
II Related work
II-A HS Object Tracker
HOT aims to discover the same objects in subsequent frames based on the initial object state and spectral information. Early work mostly used traditional methods for predicting the object position, such as histogram feature-based methods [6, 19], correlation filters-based methods [20, 21, 22], and spectral reflectance-based methods [23], which has a limited ability to deal with the complicated distractors.
Many deep learning-based methods have been recently proposed to extract robust deep features and handle complex scenes such as occlusion. It can be mainly categorized into band selection-based methods and feature fusion-based methods. The former methods leverage the RGB tracker to predict the motion of objects via the false-color images generated from HSIs. For example, Li et al. [14] proposed an ensemble learning-based method based on the false-color images generated by a band-selection method. The network produced the band weights via an auto-encoder-like structure, which is used to regroup multiple bands to several three-channel false-color images. Li et al. [24] designed an attention-based band grouping method to generate the importance weights of different bands. Chen et al. [10] proposed a spectral-spatial self-expression module to transform the HSI into false-color images by filling the modality gap. Chen et al. [11] presented a spectral awareness module to explore the band weight distributions by the multi-level band interactions. Gao et al. [25] manually selected the different bands to form the false-color images and extracted the bidirectional fused features to construct the robust features. The latter methods mainly adopt the Siamese tracker framework, comprising the spectral and RGB tracking branches to mine the cross-modality semantic cues. For example, Liu et al. [12] presented a spectral and visual feature fusion framework SiamHYPER to provide a rich semantic feature representation and obtained the accurate tracking results from the cross-modality complementary decision. Zhao et al. [26] fused features from spectral and RGB domains by the Transformer module to capture the correlation between the template images (TI) and search images (SI). Wu et al. [9] proposed a Siamese Transformer tracker to learn the heterogeneous features from different HSIs by the domain label reverse learning. Jiang et al. [27] proposed a unified network, termed as SiamCAT, to process HSIs that have varied numbers of bands. Besides, with a guided learning attention module, the spectral features can be highlighted by the spatial attention obtained from the RGB branch. Li et al. [28] designed a material-guided multiview fusion network MMF-Net, which combines the RGB, spectral, and material information to form a unified rich representation of an object.
II-B HOT Datasets
The existing HOT datasets are listed in Table I, most containing 16 or 25 bands. The detailed information is introduced as follows.
1) HOTC-2020 [6] is the first mainstream 16-band HOT dataset, consisting of 75 HSV sequences, of which 40 are used for training and 35 are used for testing, totaling 29,413 HSIs. The main shooting objects include cars and pedestrians.
2) IMEC-25 [7] consists of 80 HSV sequences for training and 55 HSV sequences for testing. The main shooting objects include cars, pedestrians, and planes, which are collected by a 25-band IMEC camera.
3) WHU-Hi-H3 [8] contains 22,753 25-band frames of annotated HSVs covering nine scenes, where 6969 frames are for training, and the remaining images are for testing. The main shooting objects include cars and pedestrians.
4) HOTC-2023 is available at the HOT competition website and consists of three sub-datasets: 15-band, 16-band, and 25-band sets, which are collected from wavelength ranges of 600-850, 460-600, and 665-960 nm, respectively. The main shooting objects include cars and pedestrians.
It can be seen that the above datasets neglect the camouflage characteristics of objects and only collect images in natural scenes, with most objects being vehicles and pedestrians, lacking the challenge of visual similarity. Although the WHU-Hi-H3 dataset considers the similarity between objects relative to the background, it does not highlight the visual confusion caused by the similarity between objects. Therefore, to facilitate research on HCOT, we construct a large-scale 25-band dataset, BihoT, which contains 41,912 frames. The main shooting objects include fruits and daily necessities.
II-C Transformer-based Object Tracker
In recent years, the Transformer has been applied in many computer vision tasks such as detection, segmentation, and classification due to its outstanding ability to model long-term dependencies. Many Transformer-based tracking methods have been proposed. Xie et al. [29] proposed a novel object-dependent feature network to embed the inter-image correlation. The non-object features can be suppressed in the self- and cross-attention schemes. Song et al. [30] designed a multi-scale cyclic shifting window attention to capture the window-level features. Mayer [31] proposed a Transformer-based model prediction module, which captures global correlations without inductive bias. Blatter et al. [32] designed an exemplar Transformer using the instance-level exemplar attention with the refined scaled dot-product calculation. Gao et al. [33] proposed an attention in attention module, which can find the consensus among all correlation vectors to enhance correlations. Cui et al. [34] presented a compact tracking framework MixFormer, which extracts and fuses features simultaneously through a mixed attention module without the need for additional fusion modules or box post-processing. Zhu et al. [16] developed a visual prompt tracking framework to achieve task-oriented multi-modality tracking. With the guidance of learned prompts, existing RGB-domain foundation models can effectively handle the downstream multi-modality tracking tasks. For HS trackers, there is currently limited research based on transformers. Gao et al. proposed a novel transformer-based tracking algorithm (Trans-HST) with a cross-band feature enhancement module to reduce the interaction information loss.
III BihoT Dataset
III-A Image Collection
We employ an IMEC XIMEA MQ022HG-IM-SM5X5-NIR camera and an EDMUND OPTICS 35mm C VIS-NIR lens to capture the high spatial and spectral resolution HS images across the wavelength range of 600-974 nm. The original image is a two-dimensional image with pixels of . After processing with the transformation algorithm, a 3D hyperspectral cube of can be obtained. To ensure the diversity of the scene, we collect data from different seasons and illuminations. The objects are real and artificial plastic fruits or vegetables in daily life to ensure the acquisition of camouflaged objects. which is shown in Fig. 3
During the image collection process, the location of the camera, exposure time, white balance, and other camera parameters are kept consistent. Besides, we found that some images do not contain the object, or the object moves out of the field of view during motion. Therefore, we carefully examine the entire dataset and filter out these noisy images.
III-B Data Annotation and Condition Challenges
The BihoT dataset has 49 annotated sequences, of which 27 are for training and 22 are for testing. Each image has an object marked by a bounding box. All bounding boxes were manually created using the Labelme toolkit. In addition, we set 9 attributes corresponding to different challenging factors in the testing set, including Background Clutter (BC), Fast Motion (FM), In-Plane Rotation (IPR), Illumination Variation (IV), Low Resolution (LR), Occlusion (OCC), Out-of-Plane Rotation (OPR), Spectral Consistency (SC), and Spectral Variation (SV). Table II lists the number of bounding boxes and challenging scenarios for each sequence. To measure the spectral feature extraction ability of the tracker, we proposed two challenging factors, Spectral Consistency (SC) and Spectral Variation (SV), compared with the existing datasets [6, 8, 7]. Specifically, in the BihoT dataset settings, there are multiple real objects and one camouflaged object in each sequence. SC means that when the real object moves, its spectral information remains unchanged relative to other real objects in the scene. When the camouflaged object moves, the spectral information relative to other real objects in the scene is different. Therefore, the scene where the camouflaged object moves is denoted as SV.
| Name | Attribute | Box Num. | Name | Attribute | Box Num. |
| Apple3 | IPR, IV, OCC, OPR, SV | 770 | Kiwifruit1 | FM, IPR, IV, OCC, OPR, SC | 649 |
| Bottle1 | OCC, SV | 631 | Kiwifruit3 | FM, IPR, IV, OCC, OPR, SV | 885 |
| Bowel1 | IPR, OCC, OPR, SV | 201 | Orange3 | BC, IV, OCC, SV | 680 |
| Bowel2 | IPR, OCC, OPR, SV | 248 | Paper | BC, SV | 900 |
| Bowel3 | OCC, OPR, SV | 102 | Tank1 | BC, SC | 820 |
| Chili1 | BC, IPR, LR, OCC, OPR, SC | 945 | Mango1 | BC, IPR, OCC, OPR, SC | 870 |
| Chili2 | BC, LR, OCC, OPR, SV | 1180 | Mango2 | BC, IPR, OCC, OPR, SV | 830 |
| Cup1 | IV, SV | 1431 | Naloong2 | IPR, OCC, SV | 1301 |
| Cup2 | IV, SV | 1666 | Lemon1 | FM, IPR, OCC, OPR, SC | 750 |
| Gear2 | LR, OCC, OPR, SC, SV | 1792 | Lemon2 | FM, IPR, OCC, OPR, SV | 620 |
| Tank2 | BC, OPR, SV | 1071 | Lemon3 | FM, IPR, IV, OCC, OPR, SV | 570 |
III-C Data Statistics
Table I lists a statistics comparison between the proposed BihoT and the existing HOT datasets. We can see that BihoT is currently the largest 25-band hot dataset, consisting of 41,912 images, with an average of 855.35 frames per video, making it a valuable dataset for developing HOT search.
To further analyze the BihoT, Fig. 4 shows the images’ motion distribution and object size changes. We first calculate the average probability density function map. The brighter area indicates the area where the object centroid frequently appears. As shown in Fig. 4(a), the bright area has a wide range and appears scattered. This indicates that the object motion in our dataset is irregular, not moving uniformly like pedestrians or vehicles on the road, which is more challenging. Then, we calculate the ratio of object pixels to images to calculate the object size. As shown in Fig. 4(b), the BihoT contains objects of different sizes, and small- or medium-sized objects have a higher proportion. On one hand, the size variation of an object is beneficial for testing the ability of the tracker to handle multi-scale objects. On the other hand, small-sized objects are more prone to occlusion and background noise, which poses a higher challenge to the robustness of the tracker. By analyzing the image information, centroid distribution, and object size changes, the characteristics of BihoT can be understood, which motivates the investigation of advanced HCOT methods.
IV Proposed Method
Fig. 5 depicts the framework overview of the proposed SPDAN, which consists of two key components. 1) Spectral Prompt-based Backbone Network: This component designs a spectral embedding module that can help establish rich spectral embedding tokens and learn joint feature representations of HS and visual by embedding spectrum as prompts in the network. Therefore, it can inherit the powerful tracking ability of the visual foundation model while solving the problems of scarcity and poor portability of HS data. 2) Distractor-aware Module: This module solves disturbance interference by first establishing a perceptual disturbance statistic that can capture frames with occlusion or background interference in the video sequence. When the SPBN prediction results are unreliable, a motion model is used to update the object’s position, which can effectively reduce the decrease in the tracking performance caused by disturbances.
IV-A Spectral Prompt-based Backbone Network
Hyperspectral camouflaged object tracking task aims to search for the object in a continuous video sequence that is highly similar to the background and other objects. Camouflaged objects usually have the same appearance (color, shape, and scale) as the real ones, except for materials, so it is difficult to utilize the visual features to distinguish a camouflaged object. To overcome this issue, a spectral prompt-based backbone network (SPBN) is proposed, incorporating spectral information into the visual features to form an effective cross-modality feature representation.
The overall network structure is shown in Fig. 5. The SPBN consists of four components, including a spectral embedding network (SEN), a cross-modality adapter (CA), a visual Transformer backbone (VTB), and a head network (HN). The input of SPBN is cross-modality image pairs, i.e., an RGB image pair (template image and search image ) and an HS image pair (template image and search image ). and denote the height and the width of a template image, and and denote the height and the width of a search image, respectively. denotes the number of bands of HS images. Different from existing methods based on Siamese networks [24, 12], our proposed SPBN can simultaneously extract features from cross-modality template images and search images and use the Transformer structure to implicitly model the relationship between template and search features, avoiding the extra relationship calculations in the Siamese network-based methods.
1) Spectral Embedding Network (SEN). To mine the spectral information and perform the cross-modality feature representations, a spectral embedding network is designed. A 3-D convolution layer and a 2-D convolution layer are jointly used for embedding spectral information, which can extract spectral information while minimizing the increase in parameters. Taking the input HS image as an example, the spectral feature extraction process for each spatial position of the th feature cube can be formulated by
where denotes the activation function and is the feature cube from the previous layer. denote the height, width, and band number of the 3-D convolutional kernel, respectively. denotes the parameters of the position , and is the bias. By doing so, the size is first changed to . D denotes the depth dimension of 3-D convolution, which equals 16 in SPDAN. Due to its calculation with sliding windows in the spectral dimension, 3-D convolution can effectively extract spectral features. After obtaining the original spectral features, They are first split and flattened into sequences of patches , and a 2-D convolution layer encodes the spatial information further to form the final token. The process can be formulated by:
where and denote the height and width of the 2-D convolutional kernel, respectively. denotes the parameters of the position . In this way, the input can be embedded as a spectral token , where is the number of tokens, is the downsampling factors, and is the feature dimensions, which equals 768 in the proposed SPDAN.
2) Cross-modality Adapter (CA). SPBN freezes the entire backbone model and only learns spectral prompts for a specific modality without training the visual Transformer backbone from scratch, enabling the model to converge quickly in the limited sample situations. With the embedding tokens, the spectral prompts can be obtained by a cross-modality adapter. As shown in Fig. 6, the initial prompts can be obtained by and . For the th layer of the Transformer, the prompts learned from the CA can be formally represented as
where , is the output feature of th Transformer layer, and denotes the th cross-modality adapter. Specifically, a linear transformation is first used to reduce the dimension of cross-modality features to reduce feature redundancy. Then, to adaptively capture the complementary modality representations, a spatial attention mechanism is used to extract the complementary modality features. Finally, the features of the two modalities are mapped back to the input dimension to generate the cross-modality prompts. The overall process can be formulated by
where denote the different linear layers, respectively. denotes the spatial attention mechanisms, and is the Hadamard product.
3) Visual Transformer Backbone (VTB). After the processing of the above two modules, the RGB and HS modalities can complement each other for prompt embedding, which can be input into a backbone model for end-to-end training. Here, we adopt the advanced visual tracking framework OSTrack [35]. Each embedded cross-modality prompt, including template and search images, is input into the Transformer encoder to extract features. It is worth noting that the parameters of this backbone model are frozen. We perform multi-stage modality embedding between each Transformer layer to extract semantic features from different stages. The forward propagation process can be formulated as
where denotes the th Transformer encode layer, including multi-head self-attention sub-network, normalization layer, and feed-forward sub-network.
4) Head Network (HN). To obtain the precise position of an object, we design a head network to predict its classification score, position offset, and size, respectively. The overall structure is shown in Fig. 5. It consists of stacked convolutional layers followed by a batch normalization layer and an activation layer. The classification branch outputs a classification response map , which is a 2-D matrix with a height of and a width of , where and denotes the response value at position of . Each point in is a probability representing the foreground scores at the corresponding position of the search region. The regression branch outputs a normalized offset map and a normalized bounding box size tensor , both of which are 3-D tensors with a depth of , a height of and a width of , where and denote the height and width of the input search image, respectively, and denotes a down-sampling ratio of the SPBN. Each point in and is a 2-D matrix indicating the offset and size of the box, respectively.
The position of the object is where the maximum value in appears, and it can be expressed as . The final bounding box can be obtained by
IV-B Distractor-aware Module
Robust joint feature representation is crucial for tracker performance. However, the robust feature is inevitably compromised under disturbances caused by occlusion or similar background distractors, leading to the drift of tracking boxes. To address this issue, we propose the distractor-aware module (DAM), which has a three-step process: perceiving, responding, and rectifying.
Step 1. Perceiving. Firstly, the tracker should perceive the occurrence of distractors. Inspired by [17], a decision confidence metric is proposed. Specifically, we calculate the average peak-to-correlation energy of the predicted score map in the classification map, which can be formulated by
where and denote the maximum and minimum values of the classification map , respectively, and denotes the number of pixels of . This DC reveals the degree of fluctuation in the response map and the confidence level of the detected objects. When this value is large, it indicates that the response map fluctuates, and only one extreme point of the response value. The predicted position is close to the true one. When encountering disturbances caused by changes in conditions, the fluctuation of the response map increases and multiple ambiguous prediction points will appear, causing a decrease in DC. Fig. 7 shows the values of DC and the attention heatmap generated by the Grad-CAM++ [39]. It can be observed that when there are disturbances such as occlusion and similar background interference (frame #0055), the prediction results become unreliable. The attention heatmap shows multiple areas with high responses, resulting in the ambiguous tracking results. In this way, the distractor can be perceived.
Step 2. Responding. When a distractor occurs, causing a decrease in DC beyond the threshold, the original tracking predicted results are misleading and may be accompanied by a mutation in positions, leading to the drift of the tracking box. Considering that, in general, the object’s motion is linear and the camera imaging frequency is high (25 FPS), the object’s position after the distractor depends on the motion state of the previous frame. Therefore, we utilize a Kalman filter, including the prediction and updating process, to predict the object position when a distractor occurs. The Kalman filter utilizes linear system state functions to estimate the optimal state through the observation data. It can be well adapted to HCOT tasks under distractors. Let represent the state vector of objects at time , where and denote the center coordinates of the object and the area and aspect ratio of the object, respectively, and denote the variation speed of the three corresponding states. In the prediction stage, the Kalman filter focuses on solving the following state transformation equations:
where denotes the prior state variable at time , denotes the posterior state variable at time , denotes the state transition matrix, denotes the state noise matrix, represents the control vector of covariance matrix , and denotes the prior estimation of the error matrix. Specifically, can be formulated by
Given , the observation variable can be obtained by
where denotes the observation noise variable, and denotes the observation matrix, which can be defined as
As for the prediction process, the posterior estimate can be obtained by correcting the prior estimation using the observation variable . The updating process can be formulated by
where is the covariance matrix, denotes the posterior estimation of state vector, is the identity matrix, denotes the Kalman gain, and denotes the prior estimation of the error matrix. To obtain the robust tracking results further, the Kalman tracker only updates the state predicted by SPBN when the DC exceeds the threshold.
Step 3. Rectifying. During the tracking process, SPBN combined with the Kalman filter can effectively handle the problem of short-term occlusion. However, when the environment or other objects are visually similar to the tracking object, especially when spectral information is similar, the DC is still high, even if there is drift in the tracking box. We call this phenomenon the “Confident Error.” At this point, the tracker chooses to trust the results of SPBN (because of its high DC), but it may lead to incorrect tracking, as shown in Fig. 8. If a perfect tracklet can be obtained, the offset of the predicted box between continuous frames will be small. In practice, there are generally only two situations where a significant offset occurs: one is due to a confident error that causes the tracking box to drift, and the other is due to the re-capture of the object when it reappears after the error result is generated by the distractors. We hope to rectify the situation with a larger offset in the first case and retain the ability of the tracker to capture the object again in the second case. Specifically, before occlusion or distractor occurs, the confidence level often gradually decreases. When confident errors occur, the DC will surge, and the tracker will incorrectly predict the position of the object, causing drift of the prediction box. To address this issue, we propose a tracking rectifying strategy that combines DC and offset. When the box drift occurs, we calculate the trend of DC within the range of the past time sliding window. The confidence level gradually decreases, indicating that confident errors will occur. To handle this situation, we directly use the object state predicted by the Kalman filter as the tracking results. By doing so, we can effectively correct the first situation without the second situation that is correct. Algorithm 1 outlines the whole tracking algorithm.
V Experiments
V-A Experiment Settings
1) Datasets. In the experiment, our proposed BihoT and two HOT datasets are selected to evaluate the proposed SPDAN. The details are described as follows.
BihoT Dataset: BihoT is captured by IMEC XIMEA MQ022HG-IM-SM5X5-NIR camera and consists of 49 HS video sequences, of which 27 are used for training and 22 are used for testing. The training part contains a total of 22,992 annotated frames, while the testing part contains 18,920 annotated frames. The dataset is annotated with 9 attributes, including Background Clutter (BC), Fast Motion (FM), In-Plane Rotation (IPR), Illumination Variation (IV), Low Resolution (LR), Occlusion (OCC), Out-of-Plane Rotation (OPR), Spectral Consistency (SC), and Spectral Variation (SV).
HOTC-2020 Dataset: HOTC-2020 was proposed at the first Hyperspectral Object Tracking Challenge in 2020. It consists of 75 HS video sequences, of which 40 are used for training, and 35 are used for testing. The training part contains a total of 12,839 annotated frames, while the testing part contains 16,574 annotated frames. The dataset is annotated with 11 attributes including Background Clutter (BC), Deformation (DEF), Fast Motion (FM), In-Plane Rotation (IPR), Illumination Variation (IV), Low Resolution (LR), Motion Blur (MB), Occlusion (OCC), Out-of-Plane Rotation (OPR), Out-of-View (OV), and Scale Variation (SV).
2) Evaluation Metrics. We choose the commonly used two evaluation metrics: the area under the curve (AUC) of the success plot and the precision rate at the threshold of 20 pixels to conduct experiments. Success plots are used to evaluate the Intersection over Union (IOU) metric between the predicted and actual bounding boxes. These plots depict the percentage of frames that meet or exceed different IOU thresholds, forming a success curve with thresholds ranging from 0 to 1. The precision rate analyzes the pixel difference between the centers of the predicted and actual bounding boxes. The precision rate curve is generated by displaying the percentage of frames where the center discrepancy is less than a specified pixel threshold, which ranges up to 50 pixels. Typically, the precision plot with the threshold of 20 is used to rank the trackers, denoted as DP_20.
3) Implementation Details. The proposed SPDAN is implemented in PyTorch and trained on a machine equipped with an NVIDIA RTX 4090 Graphics Processing Unit (GPU). The backbone network SPBN and the embedding network of RGB modality are initialized with the pre-trained parameters in [35]. The AdamW optimizer with a weight decay of 0.0001 is adopted to optimize the parameters of the network. The template and search patches are set to pixels and pixels, respectively. The network is then trained for a total of 10 epochs. and the batch size is set to 32. While the DC threshold and offset are set to 0.01 and 20, respectively.
| Methods | AUC | DP_20 | (AUC) | (DP_20) |
| B | 0.639 | 0.767 | - | - |
| B+S | 0.672 | 0.805 | +0.033 | +0.038 |
| B+S+D | 0.677 | 0.811 | +0.038 | +0.044 |
| Methods | AUC | DP_20 | (AUC) | (DP_20) |
| B | 0.693 | 0.924 | - | - |
| B+S | 0.704 | 0.938 | +0.011 | +0.014 |
| B+S+D | 0.725 | 0.968 | +0.032 | +0.044 |
| Embed Dims. | AUC | DP_20 | (AUC) | (DP_20) |
| 2 | 0.651 | 0.785 | - | - |
| 4 | 0.669 | 0.799 | +0.018 | +0.014 |
| 8 | 0.668 | 0.809 | +0.017 | +0.024 |
| 16 | 0.672 | 0.805 | +0.021 | +0.020 |
| Tracker | Publications | Backbone | UFB | FB | AUC | DP_20 |
| SiamBAN[40] | CVPR 2020 | ResNet-50 | No | No | 0.587 | 0.863 |
| SiamGAT[41] | CVPR 2021 | ResNet-50 | No | No | 0.581 | 0.827 |
| SiamCAR[42] | IJCV 2022 | ResNet-50 | No | No | 0.586 | 0.846 |
| SimTrack[43] | ECCV 2022 | ViT | No | No | 0.600 | 0.845 |
| SBT[38] | CVPR 2022 | SBT | No | No | 0.626 | 0.866 |
| SeqTrack[44] | CVPR2023 | ViT | No | No | 0.594 | 0.856 |
| BAE-Net[45] | ICIP 2020 | VITAL | Yes | No | 0.606 | 0.879 |
| SST-Net[46] | WISP 2021 | VITAL | Yes | Yes | 0.623 | 0.917 |
| SiamHYPER[12] | TIP 2022 | SiamFC | Yes | Yes | 0.678 | 0.945 |
| SEE-Net[14] | TIP 2023 | SiamFC | Yes | No | 0.654 | 0.907 |
| SiamOHOT[13] | TGRS 2023 | SiamFC | Yes | Yes | 0.634 | 0.884 |
| SPIRIT[11] | TGRS 2024 | SiamFC | Yes | Yes | 0.679 | 0.925 |
| MMF-Net[28] | TGRS 2024 | SiamFC | Yes | Yes | 0.691 | 0.932 |
| SENSE[10] | IF 2024 | SiamFC | Yes | No | 0.689 | 0.951 |
| SPDAN | Ours | ViT | Yes | Yes | 0.725 | 0.968 |
V-B Ablation Experiments
In this subsection, several ablation experiments on the proposed BihoT and HOTC-2020 datasets are conducted to evaluate the effectiveness of our methods.
1) Effectiveness of the Proposed SEN: The spectral embedding network is proposed to refine the spectral features and form a robust prompt representation. To demonstrate the effectiveness of the SEN, we first construct a prompt-based framework following [16], which is denoted as “B”. Then, the SEN is utilized to incorporate the encoded spectral information into the visual features during the token embedding stage. It is denoted as “S”. We conduct the comparative experiments between the two models on the BihoT and HOTC-2020 datasets. The experimental results are listed in Tables III and IV. It can be observed that the SEN significantly improves the performance compared to the baseline model. Specifically, on the BihoT dataset, the AUC score increases by 3.3%, and the DP_20 pixels score improves by 3.8%. On the HOTC-2020 dataset, the AUC score increases by 1.1%, and the DP_20 pixels score improves by 1.4%. This indicates that spectral information is important for the model to capture camouflaged objects,
2) Effectiveness of the Proposed DAM:
DAM is proposed to perceive the occlusion distractors and rectify the tracking results. We add the DAM to the model to show its effectiveness, which is denoted as “D”. The experimental results are listed in Tables III and IV. It can be observed that DAM further enhances the tracking performance. Specifically, DAM increases the AUC and DP on the BihoT dataset from 0.672 and 0.805 to 0.677 and 0.811, respectively. Compared to the baseline model, there are improvements of 3.8% and 4.4% in both indicators. For the HOTC-2020 dataset, DAM increases the AUC score and DP from 0.704 and 0.938 to 0.725 and 0.968, respectively. Compared with the baseline model, there are improvements of 3.2% and 4.4% in both indicators. These results validate the effectiveness of the proposed DAM, which can correct the tracking results misguided by the distractors through a proposed decision confidence metric and three-step processes.
3) The Impact of Embedding Dimension In the SEN:
To explore the impact of embedding dimension (the number of output channels by 3-D convolutional layer) on the performance of the proposed SEN, we select multiple embedding dimension settings and conduct the tracking performance experiments on the proposed BihoT. The experimental results are listed in Table V. It can be seen that the four sets of embedding parameters (i.e., 2,4,8 and 16) are selected. Among them, the setting of embedding dimension 16 achieves the best results: the AUC score reaches 0.672, and DP_20 reaches 0.805. Under optimal experimental settings, the SEN can capture more comprehensive discriminative information in the spectral dimension, providing richer spectral semantic information for the prompt module.
V-C Comparison with State-of-the-arts
Performance on the BihoT Dataset. We compare the proposed SPDAN with advanced HS and RGB trackers on the BihoT dataset, including SEE-Net [14], SiamHYPER [12], MixFormer [34], SiamBAN [40] and OSTrack [35]. SEE-Net reconstructs the HS bands through an autoencoder to learn the weight of each band. SiamHYPER proposes a Siamese tracking network that adopts independent networks for feature extraction of RGB and HS modalities. MixFormer proposes a mixed attention module to extract features, and the tracking box can be obtained without a certain head network. SiamBAN proposes a box-adaptive head network to predict the accurate object positions. OSTrack is a one-stage tracking framework, which extracts the features of the template and search images simultaneously through a visual Transformer-based backbone. The experimental results are listed in Table VI. It can be observed that the proposed SPDAN achieves the best performance in both metrics. Furthermore, the BihoT dataset contains a large number of camouflage objects, which exhibit the same appearances but different spectral information. The lack of visual features poses a significant challenge to traditional RGB trackers that rely on visual features such as color and texture. This unique setting causes substantial disparities in AUC and DP scores between the conventional RGB trackers and the proposed SPDAN. This discrepancy not only highlights the limitations of current trackers but also emphasizes the effectiveness of our methods. In addition, in the comparison methods, most HS trackers use band regrouping to generate three-channel false-color images and directly use RGB trackers to predict object positions without considering spectral information. They heavily rely on the performance of RGB trackers and are not conducive to improving the performance of HS tracking. However, SiamHYPER achieves commendable results. Unlike band regrouping-based methods, it extracts spectral features and fuses cross-modality features, which, to some extent, utilizes spectral information and alleviates the dependence on RGB trackers. It indicates the importance of spectral information for HOT, especially for camouflaged objects.
Performance on the HOTC-2020 Dataset. In this sub-section, we compare the proposed SPDAN with recent HS and RGB trackers in this sub-section. HS trackers include BAE-Net [45], SST-Net [46], SiamHYPER [12], SEE-Net [14], SiamOHOT [13], SPIRIT[11], MMF-Net [28], and SENSE [10], which are tested on HS video sequences. The RGB trackers include SiamBAN [40], SiamGAT [41], SiamCAR [42], SimTrack [43], SBT [38], and SeqTrack [44], which test on false-color video sequences. The experimental results are shown in Table VII and Figs. 10(a) and (b). Our proposed SPDAN can achieve results of 0.725 on AUC and 0.968 on DP_20, surpassing the above recent methods and proving the effectiveness of our method. The AUC of RGB trackers is mostly below 0.600, and the DP_20 is below 0.900. The performance of HS trackers is generally better than that of RGB trackers. This result proves the necessity of leveraging spectral features, as relying solely on visual features cannot handle distractors from similar objects and background clutter. Then, comparing with HS trackers, it can be observed that the use of cross-modality feature fusion performs better, indicating that fusing RGB and HS modalities can help improve performance. In addition, most existing HS trackers are based on the Siamese architecture of SiamFC [47], ignoring the advantages of the Transformer in the feature extraction. Their global relationship modeling ability is very helpful for extracting the spatial and spectral information. However, ViT-based trackers like SimTrack and SeqTrack face limitations due to their heavy reliance on the visual features and a semantic discrepancy with the robust HS features, impeding their performance enhancement. Their AUC yields 0.600 and 0.594, while the DP_20 are 0.845 and 0.856, respectively. In comparison, our SPDAN has achieved improvements of 12.5% and 13.1% in AUC and 12.3% and 11.2% in DP_20, respectively.
V-D Visualization Analysis.
To demonstrate the superiority of our method, we show the tracking results on the proposed BihoT dataset. We select four video sequences, including lemon2, apple3, chill2, and tank1. Due to the lack of RGB images collected in our dataset, we select the 1st, 9th, and 16th bands to form the false-color image. The experimental results are shown in Fig. 9. The ground truth bounding box is framed by black lines. The first column is obtained from the second frame of tracking results. The results from the consecutive frames are illustrated from the second to the fourth columns. From the first and second video sequences, we can observe that most methods struggle to accurately distinguish between camouflaged and real objects, resulting in tracking failures. In contrast, our method successfully captures both visual and spectral information of the objects, demonstrating the robustness in handling camouflaged objects. In addition, for the third video sequence, SPDAN can still perform well for smaller objects, especially when dealing with object rotation. Furthermore, The fourth sequence introduces two types of challenges related to camouflage: the first is the environmental similarity to the tank model, and the second is the visual similarity between two tanks made from different materials. It can be seen that the proposed method can effectively address these two challenges. In comparison, methods like SiamHYPER and MixFormer are misled by the camouflaged tanks during object movement. Other methods are also impacted by background clutter or camouflaged elements, further underscoring the superior performance of our method.
V-E Running Time Comparison.
To demonstrate the superiority of our method in terms of running speed, we conduct experiments on the BihoT dataset. The experimental results using frames per second (FPS) as the evaluation indicator are reported in Table VIII. All the trackers ran on a Windows machine with the configuration of an Intel(R) Core(TM) i7-13700K CPU@3.40GHz and an NVIDIA 4090 GPU. Besides, the RGB trackers ran on the three-channel false-color images, and the HS tracker ran on the HSIs. SEE-Net, SiamHYPER, and SPDAN are HS trackers, and others are RGB trackers. SPDAN achieves the best running speed among all trackers. It proves that our method has better computational efficiency compared to band regrouping or existing modality fusion methods
| Trackers | AUC | DP_20 | ||||||||||||||||
| BC | FM | IPR | IV | LR | OCC | OPR | SC | SV | BC | FM | IPR | IV | LR | OCC | OPR | SC | SV | |
| OSTrack | 0.198 | 0.271 | 0.283 | 0.341 | 0.150 | 0.363 | 0.316 | 0.222 | 0.385 | 0.271 | 0.375 | 0.352 | 0.440 | 0.179 | 0.442 | 0.387 | 0.296 | 0.474 |
| SiamBAN | 0.154 | 0.220 | 0.303 | 0.261 | 0.220 | 0.369 | 0.373 | 0.238 | 0.394 | 0.215 | 0.301 | 0.383 | 0.329 | 0.307 | 0.474 | 0.487 | 0.305 | 0.514 |
| MixFormer | 0.225 | 0.259 | 0.318 | 0.286 | 0.300 | 0.420 | 0.395 | 0.253 | 0.451 | 0.262 | 0.337 | 0.377 | 0.342 | 0.330 | 0.490 | 0.465 | 0.292 | 0.533 |
| SEE-Net | 0.191 | 0.209 | 0.236 | 0.283 | 0.249 | 0.324 | 0.317 | 0.252 | 0.353 | 0.273 | 0.310 | 0.310 | 0.385 | 0.360 | 0.429 | 0.487 | 0.346 | 0.471 |
| SiamHYPER | 0.294 | 0.131 | 0.339 | 0.344 | 0.465 | 0.420 | 0.411 | 0.261 | 0.457 | 0.410 | 0.204 | 0.444 | 0.445 | 0.628 | 0.553 | 0.563 | 0.348 | 0.617 |
| SPDAN (Ours) | 0.497 | 0.745 | 0.720 | 0.701 | 0.577 | 0.738 | 0.693 | 0.627 | 0.702 | 0.623 | 0.898 | 0.855 | 0.854 | 0.724 | 0.877 | 0.826 | 0.789 | 0.830 |
V-F Attribute-Based Evaluation.
To demonstrate the ability to handle different challenge scenarios of our proposed SPDAN, we conduct attribute-based experiments with the selected SOTA trackers on the HOTC-2020 and Bihot datasets. Fig. 10 shows the results of the HOTC-2020 dataset. For the success plot, SPDAN achieves the best results in 8 out of all 11 scenarios. This proves the effectiveness of our method to handle the complex scenarios. The SPDAN performs better in scenarios where visual features are damaged, including the loss of color, which is particularly evident in the IV and LR scenarios. In these cases, the color representation of the object is compromised, leading to a significant reduction in the availability of the visual information. Despite these challenges, our proposed SPDAN yields an AUC of 0.703, which is 9.4% higher than the second-ranked SENSE in the IV scenario, indicating the importance of spectral information in extracting robust features. Additionally, the deformation of the object presents another critical challenge in feature extraction. In scenarios such as DEF and SV, changes in the object’s shape or appearance can severely impact the ability to track it accurately, as standard visual features become unreliable. Our method achieves the best results in both metrics, proving that it can effectively extract cross-modality spectral features and has strong robustness to object deformation. Our method also demonstrates an excellent ability to deal with occluded scenes. SPDAN achieves 0.708 of the AUC score and 0.954 of the DP_20, which is 5.8% and 3.5% higher than the second-ranked method. This proves that our proposed DAM can effectively handle frequent short-term occlusion.
In addition, we also conduct attribute-based experiments on the proposed BihoT dataset. The experimental results are listed in Table IX. It is worth noting that to make more appropriate attribute comparisons, we redefine the challenge scenarios based on the characteristics of the BihoT, as detailed in Table II. Unlike the settings in HOTC-2020, we remove attributes (i.e., DEF, MB, and OV) and introduce scenarios such as SC (Spectral Consistency) and SV (Spectral Variability) to more accurately assess the utilization of spectral information. Consequently, the number of attributes evaluated is reduced from 11 to 9. It can be seen that our SPDAN exhibits superior performance in different scenarios, which further proves the effectiveness of our proposed method. Specifically, in the two spectral-related scenarios SC and SV, SPDAN achieves AUC of 0.627 and 0.702 and DP of 0.789 and 0.830. The proposed method can accurately predict the position of the object in the presence of visual similarity, spectral differences, and spectral similarity interference, demonstrating that our spectral embedding network and prompt-based learning method can accurately extract spectral features for the accurate tracking.
VI Conclusion
In this paper, we proposed a hyperspectral camouflaged object tracking task and constructed a large-scale HCOT dataset, BihoT, which contains 41,912 hyperspectral images where objects have similar appearances, diverse spectrums, and frequent occlusion. Then, we proposed a spectral prompt-based distractor-aware network (SPDAN) as a simple yet effective baseline, which combines discriminative information from the HS and RGB modalities through prompt-based learning. A spectral embedding network was designed to transform the spatial and spectral information into tokenized features. To make the tracking results more robust, we introduced a distractor-awareness module to perceive the distractor and rectify the original results. The quantitive and qualitative experimental results demonstrated the effectiveness of our method.